

********************************************************************
UARS WEEKLY STATUS REPORT - 18 FEB 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometers continue to be active under PMON control*. At the
present time, PEM energy deposition is not being generated. Recent data
configuration changes have affected PEM software causing multiple failures.
These failures are being reviewed by the PEM team. In order to correct
for new telemetry requirements, PEM will most likely need to update processing
code with a patch delivery. At this time, updated HEPS calibration parameters
will be submitted. These will influence the in-situ energy deposition
predictions, so new versions of PEM energy deposition Level 3 files will be
generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE will turn on day 049 at approximately 20:00 GMT. HALOE was last on
February 6, 2000. All engineering data were within nominal operating range
and the science data results are of excellent quality.
***********
MLS
***********
MLS operated until midday 13-feb-2000 in a mode where limb scans are performed
on the day side of the orbit at ~50N-80N latitudes, with the 205 GHz radiometer
on for measurements of ClO, HNO3 and O3.
All electronics that were on worked flawlessly, and no scan slips occurred this
observing period. Unfortunately, a large fraction of the good limb scans are
being lost due to gaps in the UARS telemetry stream.
MLS is now back in its Standby Mode.
The arctic data are being processed at JPL, and the data mapped to the 465 K
potential temperature surface (lower stratosphere) are available from the MLS
web site (http://mls.jpl.nasa.gov) by clicking on "UARS MLS Data" and then on
"UARS MLS arctic observations for 1-10 February 2000."
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
24 days WIN (5.11A) PROD (L1-L3) (520R-415C-14F=91L)
9 day SOL (4.62) PROD (L1) (9R-9C=0L)
1 day SUSIM (7.2) PROD (L1-L3) (1R-1C=0L)
The CDHF will load December 1999 current day DAAC data in February 2000.
CDHF REHOST:
- Successful processing on Rehost System.
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, SUSIM.
- Development of the Alpha 4000 Cluster Operations environments for
ingest of correlative data.
- Testing of L0 ingest and NRT continues.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 February, 2000 (Orbit 45919, GMT Day 036) through 11
February 2000 (Orbit 46023, GMT Day 042).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 037/2330 45948 Off
ACRIM 038/0040 45949 On
SUSIM 038/1322 45957 On
The beta angle ranged between the angles of 52.1 degrees to 20.6 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
3.0 degrees on 15 February 1999 (DOY 046).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 04 Febraury 2000 and 11 Febraury 2000 are provided in Appendix D.
The clock error ranged from 10.2 to -10.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 02
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 04
% ACRIM = 01
% SUSIM = 01
% HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
None to report this period.
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/021 - 2287.496849
- Transponder B: 00/031 - 2287.498836
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/041 show the observatory data loss to be 565 hours, 36
minutes, 52.393 seconds (an increase of 21 minutes 50.8 seconds since the last
report period).
The increase of 21 minutes 50.800 seconds occurred over two (2) days:
7 14 minutes 11.296 seconds on Day 035 due to an unspecified problem.
7 07 minutes 39.504 seconds on Day 036 due to an unspecified problem.
These calculations are under repair. They are being recalculated from the date
where PASS 2 operations were scheduled exclusively.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT finished the algorithms for the Safehold Telemetry Monitors (TMONUs).
Development is being delayed until OBCSTF is back online. This VAX disk pack
contains the TMON generation utilities. Once the uplink tables are completed
the FOT will deliver them to the Simulator support person for implementation
and safehold scenario generation.
MEETINGS
7 February, 2000 (DOY 038) - The FOT attended the monthly Network Operations
Directorate (NOD) meeting. Topics of discussion were the future usage of SMA
services on the TDRS H, I, and J satellites, as well as a presentation on IP
Network Operations Center. In addition, a presentation was given on the status
of the Small Coverstion Device (SCD) 5.0 delivery and implementation schedule.
These SCDs will be replacing many of the PTPs currently in place in the IP
network.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for February 15, 2000 (DOY 046). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 February 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
45922 036/0533 00-042 OBC Clk Adj 0.00 to 20.6 msec/day
45930 036/1819 00-043 NBTR NBTR NOOPS
45945 037/1821 00-044 NBTR NBTR NOOPS
45948 037/2330 00-045 HALOE Power OFF
45949 038/0040 00-046 ACRIM Power ON
45949 038/0040 00-047 OBC Clk Adj 20.6 to 0.00 msec/day
45957 038/1322 00-048 SUSIM Power ON
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 February 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
ACTEMPANA1 Y-LO Temp PCU Board 45955,7,60,61
ACTEMPANA2 Y-LO Temp Sensor Electroni 45955,7,60,61
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 February 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
45922 036/05:33 (05 Feb 00) from 0.00 to 20.6 msec/day
45949 038/00:40 (08 Feb 00) from 20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 February 2000
04 Feb 2000 - GMT Day 035 Beta = 52.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------- ---------- ----------- ----- ------- ------
1 -67.2/-89.6 -3.05/-4.62 +0.0/+0.0 16.0 0.000 0.0
2 +11.2/-28.0 +5.47/+4.51 +26.8/-16.0 25.8 1.040 21.1
3 +67.2/-5.6 +4.51/+3.23 +26.0/-16.0 25.8 1.038 20.8
11 Feb 2000 - GMT Day 042 Beta = 20.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
-- ------------- --------- ----------- ----- ------ ------
1 -67.2/-89.6 -3.05/-4.62 +0.0/+0.0 16.0 0.000 0.0
2 +33.6/-22.4 +5.79/+4.83 +31.2/-14.4 25.1 1.041 21.4
3 +89.6/-5.6 +4.51/+3.87 +30.4/-15.2 25.1 1.037 21.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 February 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/9 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512 PB
SEVNTFW00/06